# A Watermark for Black-Box Language Models

### v1

@@ 1 Introduction
It can be critical to understand whether a piece of text is generated by a large language model (LLM). For instance, one often wants to know how trustworthy a piece of text is, and those written by an LLM may be deemed untrustworthy as these models can hallucinate. This problem comes in different flavors — one may want to detect whether it was generated by a specific model or by any model. Furthermore, the detecting party may or may not have white-box access (e.g. an ability to compute log-probabilities) to the generator they wish to test against. Typically, parties that have white-box access are the owners of the model so we refer to this case as first-party detection and the counterpart as third-party detection.
The goal of watermarking is to cleverly bias the generator so that first-party detection becomes easier. Most proposed techniques do not modify the underlying LLM’s model weights or its training procedure but rather inject the watermark during autoregressive decoding at inference time. They require access to the next-token logits and inject the watermark every step of the sampling loop. This required access prevents third-party users of an LLM from applying their own watermark as proprietary APIs currently do not support this option. Supporting this functionality presents a security risk in addition to significant engineering considerations. Concretely, Carlini et al. (2024) showed that parts of a production language model can be stolen from API access that exposes logits. In this work, we propose a watermarking scheme that gives power back to the people — third-party users can watermark a language model given nothing more than the ability to sample sequences from it. Our scheme is faithful to the underlying language model and it can outperform existing white-box schemes.
@@ 2 Related Work
Watermarking outside the context of generative LLMs, which is sometimes referred to as linguistic steganography, has a long history and typically involves editing specific words from an non-watermarked text. Watermarking in the modern era of generative models is nascent — Venugopal et al. (2011) devised a scheme for machine translation, but interest in the topic grew substantially after the more recent seminal works of Kirchenbauer et al. (2023a; b) and Aaronson (2023). Many effective strategies employ some form of pseudorandom functions (PRFs) and cryptographic hashes on token $n$-grams in the input text. Kirchenbauer et al. (2023a) proposes modifying the next-token probabilities every step of decoding such that a particular subset of the vocabulary, referred to as green tokens, known only to those privy to the secret key, are made more probable. Watermarked text then is expected to have more green tokens than non-watermarked text and can be reliably detected with a statistical test. The scheme distorts the text, but with the right hyperparameters a strong watermark may be embedded with minimal degradation in text quality.
@@
Meanwhile, Aaronson (2023) proposes a clever distortion-free strategy which selects the token that is both highly probable and that achieves a high PRF value. Kuditipudi et al. (2023) applies a scheme similar in spirit to Aaronson (2023) but to improve robustness to attacks, pseudorandom numbers (PRNs) are determined by cycling through a fixed, pre-determined sequence of values called the key, rather than by $n$-grams. They compute a $p$-value using a permutation test to determine if the text was watermarked with that specific key.
@@
Lee et al. (2023) adapts Kirchenbauer et al. (2023a)’s scheme for code-generation by applying the watermark only at decoding steps that have sufficient entropy. Zhao et al. (2023) investigates a special case of Kirchenbauer et al. (2023a) for improved robustness to adversarial corruption. Fernandez et al. (2023) tests various watermarking schemes on classical NLP benchmarks and also introduces new statistical tests for detection — most notably, they suggest skipping duplicate $n$-grams during testing.
@@
Yang et al. (2023) introduces a scheme that relies on black-box access to the LLM. Their method samples from the LLM and injects the watermark by replacing specific words with synonyms. Although their approach shares the assumption of black-box LLM access, as in our work, it has limitations not present in ours: the watermarking process is restricted to words that can easily be substituted with multiple synonyms, synonym generation is powered by a BERT model (Devlin, 2018), making it computationally expensive, and the scheme is not distortion-free. Chang et al. (2024) presents PostMark, a black-box watermarking method that uses semantic embeddings to identify an input-dependent set of words. These words are then inserted into the text by an LLM after decoding. However, this approach is also not distortion-free, as the insertion of words by the LLM often results in significantly longer watermarked text.
@@
Given the weakness of many schemes to paraphrasing or word substitution attacks, some have proposed watermarking based on semantics and other features that would remain intact for common attack strategies (Liu et al., 2023b; Hou et al., 2023; Ren et al., 2023; Yoo et al., 2023). Meanwhile, others have viewed the problem through the lens of cryptography and classical complexity theory (Christ et al., 2023; Christ & Gunn, 2024). Lastly, Liu et al. (2023a) proposes an un-forgeable publicly verifiable watermark algorithm that uses two different neural networks for watermark generation and detection. Huang et al. (2023) improves the statistical tests used for detection, providing faster rates than prior work.
@@
As the deployment of watermarks to LLMs is still early and also presumably secretive, the correct threat model is still undetermined. Krishna et al. (2024) shows that paraphrasing can evade both third-party and watermarking detectors alike. Somemay posit that attacks like paraphrasing or round-trip translation are unrealistic since either they are too expensive to conduct at scale or parties in possession of a capable paraphrasing model have adequate resources to serve their own LLM. Zhang et al. (2023) show that attackers with weaker computational capabilities can successfully evade watermarks given access to a quality oracle that can evaluate whether a candidate output is a high-quality response to a prompt, and a perturbation oracle which can modify an output with a non-trivial probability of maintaining quality. Alarmingly, Gu et al. (2023) demonstrates that watermarks can be learned — an adversary can use a teacher model that employs decoder-based watermarking to train a student model to emulate the watermark. Thibaud et al. (2024) formulates tests to determine whether a black-box language model is employing watermarking, and they do not find strong evidence of watermarking among currently popular LLMs.
@@
Watermarkcdf $F$, key $K$, # cand $m$, ctx len $n$, prompt $P$, seq len $k$, LM
1:$O\leftarrow\phi$While$\neg\;\textsc{stopCond}(O)$triangleright Continue until stop token is encountered or max length reached.
@@
[…]
4:WatermarkSinglecdf $F$, key $K$, # cands $m$, ctx len $n$, prompt $P$, seq len $k$, LM
[…]
6:$(X_{1},c_{1}),\dots,(X_{j},c_{j})\leftarrow\textsc{uniqueSeqsWithCounts}((Q_{1},\dots,Q_{m}))$
[…]
8:$i^{*}\leftarrow\operatorname{argmax}_{i=1}^{j}u_{i}^{m/c_{i}}$
[…]
10:ScoreSeqscdf $F$, candidates $C$, key $K$, ctx len $n$, prefix $P$
[…]
14:$R$leftarrow ($F[s]$ for $s$ in $S$)
15:$u_{i}\leftarrow F_{|R|}\left(\sum_{j}R_{j}\right)$EndFor
[…]
17:Detectcdf $F$, tokens $X$, key $K$triangleright $p$-value-based detection.
[…]
20:$S\leftarrow\textsc{removeDuplicates}(S)$
21:$R$leftarrow ($F[s]$ for $s$ in $S$)
[…]
23:
Algorithm 2 Recursive Black-Box Watermarking
[…]
4:DetectRecursivecdf $F$, tokens $X$, keys ($K_{1},\dots,K_{t}$)
[…]
7:$y\leftarrow-2\sum_{i}{\log{P_{i}}}$triangleright Combine $p$-values using Fisher’s method.
[…]
9:
@@ 3 Algorithm
Preliminaries. We begin with some preliminaries. If $F$ is a cumulative distribution function (CDF), we let $F[s]$ (square brackets) refer to a single draw from a pseudorandom number generator (PRNG) for $F$ seeded by integer seed $s$. Let $F_{k}$ be the CDF for $\sum_{i=1}^{k}X_{i}$, where $X_{i}\overset{iid}{\sim}F$. We sometimes abuse notation and treat a distribution as its CDF (e.g. $N(0,1)(2)$ is the standard normal CDF evaluated at 2) and when the context is clear we let $-F$ be the distribution of $-X$ where $X\sim F$. Now, we detail our proposed algorithm, for which pseudocode is provided in Algorithm 1.
Let $F$ be a continuous CDF of our choosing, $P$ the input prompt, $K$ a secret integer key known only to the watermark encoder and decoder, LM a conditional language model with vocabulary $\mathcal{V}$ of size $V$, and $h$ a cryptographic hash function (e.g. SHA-256) from $\mathbb{Z}^{*}$ to $\mathbb{Z}$. Let $n$ be the number of tokens (typically 4 or 5) that serves as input to our pseudorandom function. Our PRF $g:\mathcal{V}^{*}\to\mathbb{R}$ is given by $g(w)=F[h\left(K|w\right)]$, where |||| denotes concatenation.
Watermark encoding. We sample $m$ sequences $\{Q_{1},\dots,Q_{m}\}$, each consisting of at most $k$ tokens from $\texttt{LM}\left(\;\cdot\;|\;P;\;k\right)$. Let $\{(X_{1},c_{1}),\dots,(X_{j},c_{j})\}$ be the unique sequences along with their counts from $\{Q_{i}\}$ — for example, the sequence $X_{t}$ appears $c_{t}$ times in $\{Q_{i}\}$. To score each distinct sequence $X_{t}$, we first extract its $n$-grams as $\{(X_{t,i-n-1},\dots,X_{t,i})\}_{i=1}^{|X_{t}|}$, where we allow the left endpoint to spill over only to earlier-generated tokens and not the original prompt tokens. $l$-grams are taken instead for boundary indices with only $l-1<n-1$ eligible tokens strictly left of it. We compute an integer seed for each $n$-gram $w$, as $h(K|w)$. Given a collection of seeds with their associated sequences, we deduplicate seeds across the collection. We do this by picking one instance of the seed at random and remove all remaining instances from the collection. We ensure every sequence has at least one seed by adding a random seed not already used, if necessary. For each sequence $X_{t}$, we iterate through its new seeds $S_{t}$ (order does not matter) and compute the quantity $u_{t}=F_{|S_{t}|}\left(\sum_{i=1}^{|S_{t}|}F[S_{t,i}]\right)$. Finally, we compute $i^{*}=\operatorname{argmax}_{i=1}^{j}u_{i}^{m/c_{i}}$ and choose $X_{i^{*}}$ as our watermarked sequence of length at most $k$. To generate longer texts, we run the aforementioned process autoregressively until our stopping condition, where we condition the language model on $P$ and the tokens generated thus far.
@@
Watermark detection. We treat detection as a hypothesis test, where the null $\mathcal{H}_{0}$ is that the query text was not watermarked with our scheme and secret key and the alternative $\mathcal{H}_{1}$ is that it was. While Bayesian hypothesis testing could be used, this would require choosing priors for both hypotheses, which could be challenging and a poor choice could lead to terrible predictions. Let $X$ be the query text. Akin to the encoding process, we extract $W$, the set of unique $n$-grams from $X$, permitting smaller one near the left boundary. For each $n$-gram $w_{t}$ we compute $R_{t}=F\left[h\left(K|w_{t}\right)\right]$. Under $\mathcal{H}_{0}$ (assuming that the test $n$-grams are independent), $R_{t}\overset{iid}{\sim}F$, so $\sum_{t=1}^{|W|}R_{t}\sim F_{|W|}$ giving a $p$-value $p=1-F_{|W|}\left(\sum_{t=1}^{|W|}R_{t}\right)$. Our detection score $s$ is $1-p$ (higher means more likely to be watermarked).
@@
Recursive watermarking. Since our scheme requires only a black box that samples sequences, it can be applied iteratively or recursively. Consider the following. User 1 uses User 2’s LLM service who uses User 3’s LLM service, so on so forth until User $t$. Our scheme allows User $i$ to watermark its service with its secret key $K_{i}$. Each user can then run detection using its key oblivious to whether other watermarks were embedded upstream or downstream. Furthermore, the users can cooperate in joint detection by sharing only $p$-values without revealing their secret keys. This property is valuable in the service oriented architectures of today’s technology stack.
Consider the special case that all users are actually the same entity in possession of $t$ distinct keys $\{K_{1},\dots,K_{t}\}$. Then the iterative watermarking becomes a recursive one, where $K_{i}$ is used to watermark the result of watermarking with keys $\{K_{i+1},\dots,K_{t}\}$. The entity can run Detect to get a $p$-value for each key and these $t$ $p$-values can subsequently be combined using Fisher’s method. We present this recursive scheme in Algorithm 2.

### v3

@@ 1 Introduction
It can be critical to understand whether a piece of text is generated by a large language model (LLM). For instance, one often wants to know how trustworthy a piece of text is, and those written by an LLM may be deemed untrustworthy as these models can hallucinate. The goal of watermarking is to cleverly bias the LLM so that detecting its generations becomes easier. Most proposed techniques do not modify the underlying LLM’s model weights or its training procedure but rather inject the watermark during autoregressive decoding at inference time. They require access to the next-token logits and inject the watermark every step of the sampling loop. This required access prevents third-party users of an LLM from applying their own watermark as proprietary APIs currently do not support this option. Supporting this functionality presents a security risk in addition to significant engineering considerations. Concretely, carlini2024stealing showed that parts of a production language model can be stolen from API access that exposes logits. In this work, we propose a watermarking scheme that gives power back to the people — third-party users can watermark a language model given nothing more than the ability to sample sequences from it. Our scheme is faithful to the underlying language model and it can outperform existing white-box schemes.
@@ 2 Related Work
Watermarking outside the context of generative LLMs, which is sometimes referred to as linguistic steganography, has a long history and typically involves editing specific words from an non-watermarked text. Watermarking in the modern era of generative models is nascent — venugopal2011watermarking devised a scheme for machine translation, but interest in the topic grew substantially after the more recent seminal works of kirchenbauer2023watermark; kirchenbauer2023reliability and aaronson. Many effective strategies employ some form of pseudorandom functions (PRFs) and cryptographic hashes on token $n$-grams in the input text. kirchenbauer2023watermark proposes modifying the next-token probabilities every step of decoding such that a particular subset of the vocabulary, referred to as green tokens, known only to those privy to the secret key, are made more probable. Watermarked text then is expected to have more green tokens than non-watermarked text and can be reliably detected with a statistical test. The scheme distorts the text, but with the right hyperparameters a strong watermark may be embedded with minimal degradation in text quality.
@@
Meanwhile, aaronson proposes a clever distortion-free strategy which selects the token that is both highly probable and that achieves a high PRF value. kuditipudi2023robust applies a scheme similar in spirit to aaronson but to improve robustness to attacks, pseudorandom numbers (PRNs) are determined by cycling through a fixed, pre-determined sequence of values called the key, rather than by $n$-grams. They compute a $p$-value using a permutation test to determine if the text was watermarked with that specific key.
@@
lee2023wrote adapts kirchenbauer2023watermark’s scheme for code-generation by applying the watermark only at decoding steps that have sufficient entropy. zhao2023provable investigates a special case of kirchenbauer2023watermark for improved robustness to adversarial corruption. fernandez2023three tests various watermarking schemes on classical NLP benchmarks and also introduces new statistical tests for detection — most notably, they suggest skipping duplicate $n$-grams during testing.
@@
yang2023watermarking introduces a scheme that relies on black-box access to the LLM. Their method samples from the LLM and injects the watermark by replacing specific words with synonyms. Although their approach shares the assumption of black-box LLM access, as in our work, it has limitations not present in ours: the watermarking process is restricted to words that can easily be substituted with multiple synonyms, synonym generation is powered by a BERT model (devlin2018bert), making it computationally expensive, and the scheme is not distortion-free. chang2024postmark presents PostMark, a black-box watermarking method that uses semantic embeddings to identify an input-dependent set of words. These words are then inserted into the text by an LLM after decoding. However, this approach is also not distortion-free, as the insertion of words by the LLM often results in significantly longer watermarked text.
@@
Given the weakness of many schemes to paraphrasing or word substitution attacks, some have proposed watermarking based on semantics and other features that would remain intact for common attack strategies (liu2023semantic; hou2023semstamp; ren2023robust; yoo2023robust). Meanwhile, others have viewed the problem through the lens of cryptography and classical complexity theory (christ2023undetectable; christ2024pseudorandom). Lastly, liu2023unforgeable proposes an un-forgeable publicly verifiable watermark algorithm that uses two different neural networks for watermark generation and detection. huang2023towards improves the statistical tests used for detection, providing faster rates than prior work.
@@
As the deployment of watermarks to LLMs is still early and also presumably secretive, the correct threat model is still undetermined. krishna2024paraphrasing shows that paraphrasing can evade both third-party and watermarking detectors alike. Some may posit that attacks like paraphrasing or round-trip translation are unrealistic since either they are too expensive to conduct at scale or parties in possession of a capable paraphrasing model have adequate resources to serve their own LLM. zhang2023watermarks show that attackers with weaker computational capabilities can successfully evade watermarks given access to a quality oracle that can evaluate whether a candidate output is a high-quality response to a prompt, and a perturbation oracle which can modify an output with a non-trivial probability of maintaining quality. Alarmingly, gu2023learnability demonstrates that watermarks can be learned — an adversary can use a teacher model that employs decoder-based watermarking to train a student model to emulate the watermark. thibaud2024black formulates tests to determine whether a black-box language model is employing watermarking, and they do not find strong evidence of watermarking among currently popular LLMs.
@@
1: function Watermark(cdf $F$, key $K$, # cand $m$, ctx len $n$, prompt $P$, seq len $k$, LM)
2:   $O\leftarrow\phi$
3:   while $\neg\;\textsc{stopCond}(O)$ do $\triangleright$ Continue until stop token is encountered or max length reached.
4:     $O\leftarrow O\;|\;\textsc{WatermarkSingle}(F,K,m,n,P|O,k,\texttt{LM})$
5:   return $O$
6:
7: function WatermarkSingle(cdf $F$, key $K$, # cands $m$, ctx len $n$, prompt $P$, seq len $k$, LM)
8:   $Q_{1},\dots,Q_{m}\sim\texttt{LM}\left(\;\cdot\;|\;P;\;k\right)$ $\triangleright$ Draw $m$ sequences from LM, each with at most $k$ tokens.
9:   $(X_{1},c_{1}),\dots,(X_{j},c_{j})\leftarrow\textsc{uniqueSeqsWithCounts}((Q_{1},\dots,Q_{m}))$
10:   $u_{1},\dots,u_{j}\leftarrow\textsc{ScoreSeqs}(F,(X_{1},\dots,X_{j}),K,n,P)$
11:   $i^{*}\leftarrow\operatorname{argmax}_{i=1}^{j}u_{i}^{m/c_{i}}$
12:   return $X_{i^{*}}$
13:
14: function ScoreSeqs(cdf $F$, candidates $C$, key $K$, ctx len $n$, prefix $P$)
15:   $Z\leftarrow\phi$
16:   for $X_{i}$ in $C$ do
17:     for $w$ in $\textsc{NGrams}\left(X_{i},n,P\right)$ do $\triangleright$ Don’t compute $n$-grams over original prompt.
18:      $Z\leftarrow Z\;|\;(i,\textsc{inthash}(K|w))$ $\triangleright$ Apply cryptographically secure integer hash.
19:   $Z\leftarrow\textsc{removeDuplicates}(Z)$
20:   for $i,S$ in $\textsc{sortedGroupBy}(Z)$ do $\triangleright$ Iterate through each candidate’s set of unique seeds.
21:     $R$ $\leftarrow$ ($F[s]$ for $s$ in $S$)
22:     $u_{i}\leftarrow F_{|R|}\left(\sum_{j}R_{j}\right)$
23:   return $u_{1},\dots,u_{|C|}$
24:
25: function Detect(cdf $F$, tokens $X$, key $K$, ctx len $n$) $\triangleright$ $p$-value-based detection.
26:   $S$ $\leftarrow\phi$
27:   for $w$ in $\textsc{NGrams}\left(X,n,\phi\right)$ do
28:     $S\leftarrow S\;|\;\textsc{inthash}(K|w)$
29:   $S\leftarrow\textsc{removeDuplicates}(S)$
30:   $R$ $\leftarrow$ ($F[s]$ for $s$ in $S$)
31:   return $F_{|R|}\left(\sum_{j}R_{j}\right)$ $\triangleright$  Higher score means higher likelihood of being watermarked.
32:
Algorithm 2 Recursive Black-Box Watermarking
1: function WatermarkRecursive($F$, ($K_{1},\dots,K_{t}$), $m^{t}$, $n$, $P$, $k$, LM) $\triangleright$ Sub. for WatermarkSingle.
2:   if $t=1$ then
@@
4:   else
5:     $M=\textsc{WatermarkRecursive}(F,\left(K_{2},\dots,K_{t}\right),m^{t-1},n,\;\cdot\;,\;\cdot\;,\texttt{LM})$
6:   return $\textsc{WatermarkSingle}(F,K_{1},m,n,P,k,M)$
7:
8: function DetectRecursive($F$, $X$, ($K_{1},\dots,K_{t}$), $n$)
9:   $P$ $\leftarrow\phi$
10:   for $K_{i}$ in ($K_{1},\dots,K_{t}$) do
11:     $P$ $\leftarrow$ $P\;|\;(1-\textsc{Detect}(F,X,K_{i},n))$
12:   $y\leftarrow-2\sum_{i}{\log{P_{i}}}$ $\triangleright$ Combine $p$-values using Fisher’s method.
13:   return $\chi_{2t}^{2}\left(y\right)$
14:
@@ 3 Algorithm
Preliminaries. We begin with some preliminaries. If $F$ is a cumulative distribution function (CDF), we let $F[s]$ (square brackets) refer to a single draw from a pseudorandom number generator (PRNG) for $F$ seeded by integer seed $s$. Let $F_{k}$ be the CDF for $\sum_{i=1}^{k}X_{i}$, where $X_{i}\overset{iid}{\sim}F$. We sometimes abuse notation and treat a distribution as its CDF (e.g. $N(0,1)(2)$ is the standard normal CDF evaluated at 2) and when the context is clear we let F-F be the distribution of X-X where $X\sim F$. Now, we detail our proposed algorithm, for which pseudocode is provided in Algorithm 1. A Python implementation and example usage are presented in the Appendix.
Let $F$ be a continuous CDF of our choosing, $P$ the input prompt, $K$ a secret integer key known only to the watermark encoder and decoder, LM a conditional language model with vocabulary $\mathcal{V}$ of size $V$, and $h$ a cryptographic hash function (e.g. SHA-256) from $\mathbb{Z}^{*}$ to $\mathbb{Z}$. Let $n$ be the number of tokens (typically 4 or 5) that serves as input to our pseudorandom function. Our PRF $g:\mathcal{V}^{*}\to\mathbb{R}$ is given by $g(w)=F[h\left(K|w\right)]$, where $|$ denotes concatenation.
Watermark encoding. We sample $m$ sequences $\{Q_{1},\dots,Q_{m}\}$, each consisting of at most $k$ tokens from $\texttt{LM}\left(\;\cdot\;|\;P;\;k\right)$. Let $\{(X_{1},c_{1}),\dots,(X_{j},c_{j})\}$ be the unique sequences along with their counts from $\{Q_{i}\}$ — for example, the sequence $X_{t}$ appears $c_{t}$ times in $\{Q_{i}\}$. To score each distinct sequence $X_{t}$, we first extract its $n$-grams as $\left\{(X_{t,i-n-1},\dots,X_{t,i})\right\}_{i=1}^{|X_{t}|}$, where we allow the left endpoint to spill over only to earlier-generated tokens and not the original prompt tokens. $l$-grams are taken instead for boundary indices with only l1<n1l-1<n-1 eligible tokens strictly left of it. We compute an integer seed for each $n$-gram $w$, as $h(K|w)$. Given a collection of seeds with their associated sequences, we deduplicate seeds across the collection. We do this by picking one instance of the seed at random and remove all remaining instances from the collection. We ensure every sequence has at least one seed by adding a random seed not already used, if necessary. For each sequence $X_{t}$, we iterate through its new seeds $S_{t}$ (order does not matter) and compute the quantity $u_{t}=F_{|S_{t}|}\left(\sum_{i=1}^{|S_{t}|}F[S_{t,i}]\right)$. Finally, we compute $i^{*}=\operatorname{argmax}_{i=1}^{j}u_{i}^{m/c_{i}}$ and choose $X_{i^{*}}$ as our watermarked sequence of length at most $k$. To generate longer texts, we run the aforementioned process autoregressively until our stopping condition, where we condition the language model on $P$ and the tokens generated thus far.
@@
Watermark detection. We treat detection as a hypothesis test, where the null $\mathcal{H}_{0}$ is that the query text was not watermarked with our scheme and secret key and the alternative $\mathcal{H}_{1}$ is that it was. While Bayesian hypothesis testing could be used, this would require choosing priors for both hypotheses, which could be challenging and a poor choice could lead to terrible predictions. Let $X$ be the query text. Akin to the encoding process, we extract $W$, the set of unique $n$-grams from $X$, permitting smaller one near the left boundary. For each $n$-gram $w_{t}$ we compute $R_{t}=F\left[h\left(K|w_{t}\right)\right]$. Under $\mathcal{H}_{0}$ (assuming that the test $n$-grams are independent), $R_{t}\overset{iid}{\sim}F$, so $\sum_{t=1}^{|W|}R_{t}\sim F_{|W|}$ giving a $p$-value $p=1-F_{|W|}\left(\sum_{t=1}^{|W|}R_{t}\right)$. Our detection score $s$ is 1p1-p (higher means more likely to be watermarked).
@@
Recursive watermarking. Since our scheme requires only a black box that samples sequences, it can be applied iteratively or recursively. Consider the following. User 1 uses User 2’s LLM service who uses User 3’s LLM service, so on so forth until User $t$. Our scheme allows User $i$ to watermark its service with its secret key $K_{i}$. Each user can then run detection using its key oblivious to whether other watermarks were embedded upstream or downstream. Furthermore, the users can cooperate in joint detection by sharing only $p$-values without revealing their secret keys.
Consider the special case that all users are actually the same entity in possession of $t$ distinct keys $(K_{1},\dots,K_{t})$. Then the iterative watermarking becomes a recursive one, where $K_{i}$ is used to watermark the result of watermarking with keys $(K_{i+1},\dots,K_{t})$. The entity can run Detect to get a $p$-value for each key and these $t$ $p$-values can subsequently be combined using Fisher’s method. We present this recursive scheme in Algorithm 2.
@@
Relation to similar work.
With our algorithm presented, we now discuss its relation to two similar methods whose treatment we had deferred.
SynthID (dathathri2024scalable) is a white-box scheme which resembles the recursive variant we propose. Specifically, when the sequence length is 1 and each watermarking party samples 2 sequences, then our encoding procedure overlaps with theirs. Both processes can be visualized via a binary tree where nodes represent tokens. At the very bottom, leaf nodes represent tokens sampled i.i.d. from the next-token distribution. The token for a non-leaf node is decided by applying a specific PRF determined by the node’s level to the two children tokens. Scoring, or decoding, differs however in that while they propose training a Bayesian detector or zero-shot detection by combining PRF values for the unique $n$-grams across levels of the tree (equivalently, virtual watermarking parties) using a weighted average (or corresponding $z$-scores) that accounts for entropy differences across depths, we suggest combining $p$-values using Fisher’s method with equal contributions across the layers. Furthermore, they suggest using a Bernoulli distribution instead of the standard uniform and consider much deeper trees (e.g. a depth of 30) than the ones we do in this work.
Most similar to our work is the WaterMax algorithm of giboulot2024watermax. At a high level, their scheme is very similar to ours: i.i.d. responses are sampled from the LLM for a specific prompt, pseudorandom values are determined for each token in each sequence based on $n$-grams and an aggregate score is computed for each sequence. The sequence with the largest aggregate score is returned as the watermarked text. However, very crucially, their algorithm is not distortion-free, as ours is, and their theory does not capture the critical role of entropy on detectability. Meanwhile, we are able to formally and empirically quantify the dependence on entropy on performance and formulate provably optimal tests. Furthermore, we propose a recursive variant of the algorithm and also carefully study the effect of the distribution used for the PRF.